MICQ(7) MICQ(7) NAME mICQ−interactivecommands DESCRIPTION ThismanualpagedescribesthecommandsthatcanbeusedwithinmICQ.Thispageholdsthe sameinformationastheonlinehelp,butmorepreciseandlonger.Notations:squarebrackets([]) denoteoptionalarguments,whileanglebrackets(<>)denotearguments(inoppositetokeywords).Specifically,<uin>denotesanarbitraryUIN,while<contacts>denotesacommaseparatedlistofcontacts,thatis,ofUINsandnicknames.Ifitisthelastargument,thecontactsmay alsobeseparatedbyspaces. COMMANDSCONCERNINGMESSAGESENDING msg<contacts>[<message>] Sendamessagetoallcontacts.Withoutmessage,orwithmessageendinginabackslash (\),multi−linemessagemodeisenabled.Tosendthemulti−linemessage,putaperiod (.)onalineofitsown.Tocancelthemessage,putahashsign(#)onalineofitsown. Outgoingandincomingmessagesaremarkeddifferentlydependingofthetransportused. <<<<<<and>>>>>>markmessagessentviaadirectconnection.<<%<<and>>%>>mark messagessentviaadirectconnectionthatwereencrypted(requiresGnuTLSorOpenSSL enabledatcompiletime).<<<<and>>>>markicq8(alsocalledtype−2)messages. Lastly,<<<and>>>markicq5messages,whichcanbesocalledtype−1,type−4or offlinemessage.Thesekindofmessagesarenotacknowledged. Amessageisfirsttriedtobesentviaapre−existingdirectconnection;ifnoneisestablished,oneisinitiatedinthebackgroundforfurthermessages.Ifnodirectconnectionis open,orifthemessageisnotacknowledgedaftersometime,theconnectionwillbefailed andthenexttransporttried:sendingastype−2message.Thisstepwillbeskippedifthe peerdoesnotsettherequiredcapabilitiestosignalitsabilitytoreceivethem.Ifthisis skipped,timesoutorreturnsanerror,e.g.thepeerhasgoneofflineinthetimebetween, themessageissentasanordinarytype−4message. Ifthepeersignalsitsabilitytoacceptthem,messagesaresentUTF−8encodedand appropriatelytagged.Otherwise,themessageissentencodedinthepeer´ sconfigured encoding,or,ifunset,thedefaultassumedremoteencoding.type−1,type−4andoffline messagescannotbetaggedwiththeirencoding,sosuccessfultransmissionof non−USASCII−charactersreliesoncorrectconfigurationandheuristics. a[<message>] Sendamessagetothelastpersonyousentamessageto.Auto-expandingaliasfor msg%a. r[<message>] Replytolastmessagereceived.Auto-expandingaliasformsg%r. url<contacts><url><message> Sendamessageregardinganurltoallcontacts. sms[<nick>][<cell>]<message> SendanSMSmessagemessagetocellphonenumbercell,whichlookslike+<country code><number>,ortothecellphonenumberofnick,whichwillbesettocellifunset. Youmaynotspecifycellifnickalreadyhasacellphonenumbergiven.Invalidcellphone numberswillberemovedfromthenick´ smetadata. chat<contacts>[<message>] Sendsamessagemessagetocontactscontactsjustasthemsgcommand.Afterthemessageissent(orcancelled),multi-linemodeisenteredandmessagessentuntilanempty messageiscancelled.Cancellinganon-emptymessagewillonlycancelthemessageitself, andmoremessagesinmulti-linemodecanbesent. mICQ 1 MICQ(7) MICQ(7) getauto[auto|away|na|dnd|occ|ffc][<contacts>] Requestsautomessagefromcontactsforgivenstatus,whereawayisforaway,nafornot available,dndfordonotdisturb,occforoccupiedandffcforfreeforchat.Ifautoor nothingisgiven,theautomessageforthecontactsisfetchedfortheirrespectivestatus. Contactsnotinanyofthosestatuswillbeskipped.Newfor0.4.10. auth[req|grant|deny|add]<contacts> Grantordenyallcontactstoaddyoutotheirlist,requestfromallcontactstoallowyou toaddthemtoyourcontactlist,ortellcontactsthatyouaddedthemtoyourcontact list. resend<contacts> Resendthelastmessagetomorecontacts. last[<contacts>] Showthelastmessagereceivedfromcontact,orfromeveryonewhoalreadysentamessage. h<contact>[<last>[<count>]] history<contact>[<last>[<count>]] Showmessagesfrom<contact>fromthelogfiles.Startatthelastlatestmessage,orthe -lastmessageiflastisnegative,defaultingtothe20iethlastmessage,anddisplaycount messages,defaultingto20ifnotgiven. historyd<contact><date>[<count>] Showmessagesfrom<contact>fromthelogfiles.Startatdatedate,whichmustbegiven inISO8601format,eitherYYYY−MM−DDorYYYY−MM−DDTHH:MM:SS,anddisplaycountmessages,defaultingto20ifnotgiven. find<contact><pattern> Searchthepatternpatterncase−insensitiveinthelogfileforcontact. finds<contact><pattern> Searchthepatternpatterncase−sensitiveinthelogfileforcontact. tabsShowsofalistofusersyoucantabthrough. COMMANDSCONCERNINGYOURSTATUS loginAbuilt−inshortcutforconnlogin.Opensthefirstserverconnection. onlineChangestatusto"online". away[<message>] Changestatusto"away".Optionallysetautoresponsetemporarelyforthisstatusto message. na[<message>] Changestatusto"notavailable".Optionallysetautoresponsetemporarelyforthisstatus tomessage. occ[<message>] Changestatusto"occupied".Optionallysetautoresponsetemporarelyforthisstatusto message. dnd[<message>] Changestatusto"donotdisturb".Optionallysetautoresponsetemporarelyforthisstatustomessage. ffc[<message>] Changestatusto"freeforchat".Optionallysetautoresponsetemporarelyforthisstatus tomessage. invChangestatusto"invisible". mICQ 2 MICQ(7) MICQ(7) change[<number>[<message>]] Changesyourstatustonumber.Withoutanumberitlistssomeavailablemodes.Optionallysetautoresponseforthisstatustomessage. COMMANDSCONCERNINGFINDINGANDSEEINGOTHERUSERS f<contacts> finger<contacts> Showallwhitepageinformationforcontacts<contacts>. ss<contacts> Showallsavedwhitepageinformationforcontacts<contacts>. iListallthepeopleonyourignorelist. s[<contacts>] Showyourcurrentstatus,orofallgivencontactsindetail,includingallaliases. e,ee,eg,eeg,ev,eev,egv,eegv,w,ww,wg,wwg,wv,wwv,wgv,wwgv Listpartsofyourcontactlist.Thefollowingflagsaredisplayedinthefirstcolumn: +Thisentryisactuallynorealcontact,butanaliasoftheprecedingone.Only withww. #Thisentryisnotonyourcontactlist,buttheUINwasusedatsometime.Only withwandww. *Thiscontactwillseeyourstatusevenifyou´ reinvisible. −Thiscontactwon´ tseeyouatall. ˆ Thiscontactisignored:nomessages,nostatuschanges. Theww*andee*groupofcommandsdisplayanothercolumnfordirectconnectionflags. Ifthefirstcolumnwouldbeempty,itwillbedisplayedinthefirstcontactforallofthose commandsunlessit´ sˆ . &Adirectconnectiontothiscontactisestablished. |Adirectconnectiontothiscontacthasfailed. :Adirectconnectioniscurrentlytriedtobeestablished. ˆ Nodirectconnectionopened,butIPaddressandportnumberareknown. Anothercolumnisaddedfortheww*andee*groupofcommandstoshowthestatus withregardstotheserversidecontactlist,asfarasitisknown: SThecontactisintheserversidecontactlist,andshouldbethere. sThecontactisintheserversidecontactlist,butshouldnotbethere. .Thecontactisnotintheserversidecontactlist,butshouldbethere.Thisusuallymeans"authorizationrequired". ´´ Thecontactisnotintheserversidecontactlist,andisnotwantedthere. e[<group>] Listallpeoplewhichareonlineinyourcontactlist.Alsoprintyourownstatus. ee[<group>] Listallpeoplewhichareonlineinyourcontactlistwithmoredetails. eg,eeg Asabove,butsortbygroups.Newfor0.4.10. w,ww,wg,wwg Asabove,butalsoincludeofflinecontacts. mICQ 3 MICQ(7) MICQ(7) ev,eev,egv,eegv,wv,wwv,wgv,wwgv Asabove,butalsoincludehiddencontacts(withtheshadowoptioneffectivelyset).New for0.5. ewideListallpeoplewhichareonlineinyourcontactlistinascreenwideformat. wideListallpeopleinyourcontactlistinascreenwideformat. search search<em@il> search<contact> search<first><last> Searchforauserwithem@ilastheiremailaddress,withnickastheirnick(which mustn´ tcontainan@),orwithfirstastheirfirstandlastastheirlastname.Ifnoargumentisgiven,askfornickname,firstname,lastname,emailaddressandawholebunch ofotherdatatosearchfor. rand[<number>] Findarandomuserininterestgroupnumber.Ifnumberisnotgiven,listallinterest groups. COMMANDSCONCERNINGYOURCONTACTLIST add<uin><nickname> Adduintoyourcontactlistasnickname. Note:youneedtosavetomakethispersistent. add<contact><alias> addalias<contact><alias> Addaliasaliasfornickname.New(addalias)for0.4.10. Note:youneedtosavetomakethispersistent. add[<group>[<contacts>] addgroup<group>[<contacts>] Addallcontactsincontactstothecontactgroupgroup,whichiscreatedifitdoesn´ t existandthecommandisgivenasaddgroup.Newfor0.4.10. Note:youneedtosavetomakethispersistent. rem[all]<contacts> remalias[all]<contacts> Removethealiasesincontacts.Ifanynickisthelastaliasforthiscontact,orifallis given,removethecontactcompletely. New(remalias)for0.4.10.Note:youneedtosavetomakethispersistent. remall<group> rem<group><contacts> remgroupall<group> remgroup<group><contacts> Removeallcontactsincontactsfromthecontactgroupgroup,orremovethecontact groupcompletelyifallisgiven.Newfor0.4.10.Note:youneedtosavetomakethis persistent. togig<contacts> Toggleswhethercontact´ smessagesandstatuschangesareignored. toginv<contacts> Toggleswhetheryouwillbehiddenforcontacts. mICQ 4 MICQ(7) MICQ(7) togvis<contacts> Toggleswhethercontactscanseeyouevenifyouareinvisible. COMMANDSCONCERNINGYOURICQACCOUNT pass<password> Changesyouricqpasswordtopassword. Note:yourpasswordmaynotstartwitho´ (thebyte0xf3). Note:youneedtosavetomakethispersistentincaseyousaveyourpasswordinyour ˜ /.micq/micqrcfile(seemicqrc(7)),oryouwillgetamismatchedpasswordonyournext login. update Updatesyourbasicuserinformation(email,nickname,etc.). otherUpdatesotheruserinformationlikeageandsex. aboutUpdatesyouraboutuserinformation. setr[<number>] Setsyourrandomusergrouptonumber.Withoutargument,listspossibleinterest groups. reg<password> Createsanewuseraccountwithpasswordpassword. COMMANDSFORCONFIGURINGTHEMICQCLIENT verbose[<level>] SetverbosityofmICQtolevel,orshowthecurrentverbosity.Theverbositylevelisaset offlags: 8Showprotocolerrors. 16Showcreationanddeletionofpackets. 32Showaddedandremovedqueueevents. 64Showcreatedanddeletedconnections. 256Showv6packetsandtheirdata. 4096Showv7/v8packets. 8192Showthehexdumpofv7/v8packets. 16284Savethosepacketstodisc. 65536Showpeer−to−peerpackets. 131072 Showthehexdumpofpeer−to−peerpackets. 262144 Savethosepacketstodisc. 2097152 Showpeer−to−peerhandshake. 4194304 ShowI/Oconnectionsetup. However,anynon−zeroverboselevelwillcausemoreinformationtobeprinted,andany levelgreaterthan1evenmore. clearClearthescreen. sound[on|off|event] Togglesoundonoroff,orcalltheeventscriptforbeeps.Newsyntaxfor0.4.10. mICQ 5 MICQ(7) MICQ(7) autoaway[<timeout>|on|off] Setthetimeouttotimeout,orswitchautoawayonusingthelastusedordefaultauto awaytime,orswitchautoawayoffsavingthecurrentusedautoawaytime.Atimeoutof 0isequivalenttooff.Ifnoargumentisgiven,thecurrentsettingisdisplayed. alias[auto[expand]][<alias>[<expansion>]] Makeanewaliasnamedaliaswhichissubstitutedwithexpansion.Ifnoaliasisgiven, listallcurrentaliases.Ifnoexpansionisgiven,listthecurrentalias.Otherwise,createa newalias.Iftheautoorautoexpandkeywordisgiven,thenewaliasisauto-expanding, thatis,itisimmediatelyexpandedwhenthespaceorreturnkeyispresseddirectlyafter enteringthealias. Ifthestring%sispresentinexpansion,eachoccurenceisreplacedwiththegivenargumentswhenthealiasisinvoked,otherwisetheywillbeappended.Ifthestring%ris presentinexpansion,eachoccurenceisreplacedbythenicknameor,ifthenickisnot available,theUIN,ofthecontactthatsentthelastmessagetoyou,ortheemptystringif thereisnone.Ifthestring%aispresentinexpansion,eachoccurenceisreplacedbythe nicknameor,ifthenickisnotavailable,theUIN,ofthecontactthatyousentthelast messageto,ortheemptystringifthereisnone. Newfor0.4.10.Auto-expandingaliasesnewfor0.5.0.4. unalias<alias> Deletethealiasnamedalias.Newfor0.4.10. lang[<lang>|<nr>]... Switchtothelanguageandencodinglang,givenbyisocode,oroutputtranslationstring numbernr.ToselectthedefaultlanguagegivenbytheenvironmentvariablesLANG, LCALLandLCMESSAGES,use!,autoordefault.Toselectnotranslationatall, use.,noneorunload.Togetdebugginginformation,prefixitwithdebug. Thefilesfrom˜ /.micq/i18n/*.i18ntakeprecedenceovertheglobalones,thefilesfrom BASE/i18n/*.i18btakeprecedenceoverthosefrom˜ /.micq/i18n/*.i18n,whilethe "funny"textstakeprecedenceoverthestandardones. Youmayspecifyasmanyargumentsasyouwish,e.g.toprintstringnumber117inGermanandthenswitchbacktodefaultlanguage,use"transde117default". uptime ShowthetimemICQhasbeenrunning,plussomestatistics. set<option><value> Setoptionoptiontoeitheronoroff,oranoptionspecificsetofvalues.optioncanbe: color:usecolors, delbs:interpretdeletecharactersasbackspace, funny: usefunnymessages, auto:sendauto−replieswhenyou´ ree.g.away, uinprompt: havethelastnickintheprompt, autosave: automaticallysavethemicqrc, autofinger: automaticallyfingernewUINs, mICQ 6 MICQ(7) MICQ(7) linebreak: setthestyleforline−breakingmessagestosimple,tosimplyprintthemessage afterthenickandwrap,break,toaddalinebreakbeforeprintingeachincoming message,indent,toindentalllinesofamessagetothelevelafterthenick,or smart,toaddalinebreakonlyifthemessagewouldn´ tfitontheline; Newoptionsfor0.4.10.Optionshermit,log,logonoff,silent,tabsremovedfor0.5;usethe optcommandforthem. opt[<contact>|<contactgroup>|connection|global[<option>[<value>]]] Setoptionoptionforcontactcontact,contactgroupcontactgroup,forthecurrentconnectionorgloballytothevaluevalue,ordisplaythecurrentvalue,ordisplayalloptions andtheirsettings. Whenresolvingoptionsforcontacts,theoptionforthecontactitselfwillbeused,ifset; otherwise,ifthecontactbelongstoagroup,theoptionforthatgroup,ifset;otherwise,if thecontactbelongstoaserverconnection(i.e.isnotyetremoved),theoptionforthe serverconnection;otherwisetheglobaloption,ifset.Otherwise,0forintegeroptions, FALSEforbooleanoptions,thecolorstringfor"nocolor"unlessthiscolorstringis requested,ortheemptystringforstringoptions. Whenresolvingoptionsforcontactgroups,thesameisdoneexceptforthefirststep; optionsforconnectionsaretakenfromtheconnection,thenglobally;andglobaloptions, ofcourse,onlyglobally. Thefollowingoptionscurrentlyexist,withtypeandapplicability: colorschemeintegerglobal Thenumberofthecolorschemetouse.Setto0todisableandusethecolors definedbelow. colornonecolorglobal Thecolorstringtousefornormaltext.Thekeywordsblack,red,green,yellow,blue,magenta,cyan,white,none,andboldareunderstoodandreplaced bytheirANSIcontrolsequences.Notethatanycoloralreadyincludesnone,so specifyboldalwaysafterthecolor.Anyothertextisinterpretedascontrolcharacterstooutputverbatim. colorservercolorglobal Thecolorstringtouseforserverstuff. colorclientcolorglobal Thecolorstringtouseforclientstuff. colorinvcharcolorglobal Thecolorstringtousewhendisplayingbyte−sequencesinvalidforthegiven encoding. colorerrorcolorglobal Thecolorstringtouseforerrors. colordebugcolorglobal Thecolorstringtousefordebugoutput. colorquotecolorglobal Thecolorstringtouseforquotedtext. webawarebooleanconnection Whetherthecurrentstatusshouldbevisibleontheweb. hideipbooleanconnection Whethertohidethe(local)IPaddress. mICQ 7 MICQ(7) MICQ(7) dcauthbooleanconnection Whetherauthorizedcontactscanseethe(remote)IPaddress. dccontbooleanconnection Whethercontactscanseethe(remote)IPaddress. obeysblbooleanconnection Whethertoloadtheserver−basedcontactlistuponlogin.Thiswillenforce requiredauthorizations,whichwillbeignoredforlocalcontactlists. ignorebooleancontact Whethertoignoreallmessagesfromthiscontact. hidefrombooleancontact Whethertobeinvisibleforthiscontact. intimatebooleancontact Whethertobeallwaysvisibleforthiscontact. logonoffbooleancontact Whethertologwhenthiscontactsgoesonlineoroffline. logchangebooleancontact Whethertologstatuschangesforthiscontact. logmessbooleancontact Whethertologmessagesforthiscontact. showonoffbooleancontact Whethertoshowwhenthiscontactgoesonlineoroffline. showchangebooleancontact Whethertoshowstatuschangesforthiscontact. autoautobooleancontact Whethertoautomaticallyrequestautomaticmessagesonstatuschangesforthis contact. hideackbooleancontact Whethertohideacknowledgesformessagestothiscontact. Note:usingthisisareallybadidea,asyou’llmisswhenmessagesareresentwithoutacknowledge. wantsblbooleancontact Whetheryouwantthiscontacttobeontheserversidecontactlist. peekmebooleancontact Whetheryouwantthiscontacttobepeekedwhenusingthepeekallorpeekall command. shadowbooleancontact Whetheryouwanttohidethiscontactincontactlistdisplaysaslongasnoneof the*vcommandsisused. localbooleancontact Whetherthiscontactisnotarealcontact,butjustanickforanumber. encodingstringcontact Theencodingtoassumeformessageswhoseencodingisundefined,inparticular metadataontheserver. tabspoolintegercontact Whethertoaddacontacttothetabslistuponstartup. mICQ 8 MICQ(7) MICQ(7) autoawaystringcontact Theautoawaystringforstatusawaytoreturntothiscontact. autonastringcontact Theautoawaystringforstatusnatoreturntothiscontact. autooccstringcontact Theautoawaystringforstatusocctoreturntothiscontact. autodndstringcontact Theautoawaystringforstatusdndtoreturntothiscontact. autoffcstringcontact Theautoawaystringforstatusffctoreturntothiscontact. colormessagestringcontact Thecolorstringtouseformessagesofthiscontact. colorsentstringcontact Thecolorstringtouseforthenickwhenmessagesaresent. colorackstringcontact Thecolorstringtouseforthenickwhenmessagesareacknowledged. colorincomingstringcontact Thecolorstringtouseforthenickwhenmessagesarereceived. colorcontactstringcontact Thecolorstringtouseforthenickotherwise. Newfor0.5. optcontact<contact>[<option>[<value>]] Setordisplayoptionoptionoralloptionsofcontactcontact.Newfor0.5. optgroup<group>[<option>[<value>]] Setordisplayoptionoptionoralloptionsofcontactgroupgroup.Newfor0.5. optconnection[<option>[<value>]] Setordisplayconnectionoptionoption.Newfor0.5. optglobal[<option>[<value>]] Setordisplayglobaloptionoption.Newfor0.5. saveSavescurrentsettingsintheconfigurationfile,whichis˜ /.micq/micqrc,unlessspecified otherwise.Beware,itwillclobberanycommentsinthisfile. q[<msg>] quit[<msg>] exit[<msg>] QuitmICQ.Ifgiven,sendthemessagemsgtoallcontactspreviouslymessagedthatare stillonlineandonthecontactlist. Note:mICQwillnotwaitforanyacknowledges-ifthemessageisnotreceivedbythe contactforwhateverreason,itwillnotberesent. Aliasesquitandexitnewfor0.4.10. x[<msg>] QuitmICQwithoutsaving.Ifgiven,sendthemessagemsgtoallcontactspreviouslymessagedthatarestillonlineandonthecontactlist. Note:mICQwillnotwaitforanyacknowledges-ifthemessageisnotreceivedbythe contactforwhateverreason,itwillnotberesent. Aliasesquitandexitnewfor0.4.10.Newfor0.5. mICQ 9 MICQ(7) MICQ(7) COMMANDSCONCERNINGADVANCEDFEATURES meta[show|load|save|set|get|rget]<contacts> Handlemetadataofcontacts.Thefollowingsubcommandsareknown: showShowsthemetadataofallcontactsgiven. loadLoadsthemetadataforallcontactsgivenfromdiscandshowsthem. saveSavesthemetadataforallcontactsgiventodisc. setUploadsyourmetadatatotheserver. getGetsthemetadataforallcontactsfromtheserverandshowsthem. getrGetsthemetadataforthecontactthelastmessagewasreceivedfromfromthe serverandshowsit. Newfor0.4.10. file[...] Aliasforpeerfile.Newfor0.4.10. accept[...] Aliasforpeeraccept.Newfor0.4.10.4. peer<command><uin|nick> OperatecommandcommandonusergivenbyUINuinornicknamenick. openOpenapeertopeerconnectionoverTCPtotheuser. closeCloseandresetapeertopeerconnectiontotheuser. offSwitchofftryingtoestablishsuchaconnectionforsendingmessagesuntilitis explicitlyopenedorreset. file<file><description> Sendfilefilewithdescriptiondescription. files[<file><as>]...<description> Sendfilestotheuser.Theremaybearbitrarilymanypairsofaphysicalfilename fileandthenametobepresentedtothepeer,as.Ifasis´ /´ ,thefilename withoutthepathissent,andifitis´ .´ thesamenameissent. accept[<contact>][<id>] AcceptanincomingfilerequestfromcontactwiththegivenIDid.Ifthereis onlyonependingincomingfilerequest,thecontactandidargumentsmaybe omitted,otherwiseitisundefinedwhichfilerequestisansweredwhenseveral match. deny[<contact>][<id>][<reason>] DenyanincomingfilerequestfromcontactwiththegivenIDidwiththereason reason. login[...] Aliasforconnlogin.Newfor0.4.10.4. conn[<command><nr>] Listallconnections,oroperatecommandonconnectionnr. openOpenthegiven,orthefirstserverconnection. loginOpenthegiven,orthefirstserverconnection. closeClosethegivenconnection.Temporaryconnectionswillgetremovedbythis. remove Closeandremovegiven(temporary)connection. mICQ 10 MICQ(7) MICQ(7) selectSelectthegivenserverconnectionasthecurrentone.nrmaybetheconnection numberortheUINusedfortheconnection. contact[<command>] Handletheserversidecontactlist: showDownloadtheserversidecontactlistandjustshowit. diffDownloadtheserversidecontactlistandshowonlycontacts(uin/nickpairs)that arenotinthelocalcontactlist. addDownloadtheserversidecontactlistandaddallcontactstothelocalone. upload Trytouploadlocalcontactstotheserversidecontactlist. download Downloadcontactsfromtheserversidecontactlist,butavoidmodifyingalready existingcontacts. import Downloadcontactsfromtheserversidecontactlist,modifyingexistingcontacts. Note:Thiswillalwaystrytouploadonlythosecontactswiththewantsblcontactoption.Doaoptglobalwantsblonifyouwanttouploadallofyourcontacts. peek[<contacts>][all][<contacts>] Checkforeachcontactwhetheritisactuallyonlineornot.Thespecialkeywordallwill checkuponallcontactswiththepeekmeoptioneffectivelyset. Note:ThisabusesabugintheICQprotocoltofigurethisoutandthuscanstopworking atanytime.Noadditionalinformationexceptonlineorofflinecanbefoundoutthisway. peek2<contacts> Builtinaliasforgetautoawaycontacts.Canbeusedtodetectpresenceofusersofsome clients,butthisitselfcanbedetectedbytheotheruser’sclient.mICQsince0.5willnot bedetectedbythis,butwilldetecttriesfromotherclients. peekall[<contacts>] Builtinaliasforpeek<contacts>all. as<nr|uin><cmd> ExecutemICQcommandcmdasiftheconnectionnumbernrorfortheUINuinwould bethecurrentone. COMMANDSFORSCRIPTING tclscript<file> Executethetclscriptfile,whichmaybegivenrelativetothemICQbasedirectory. tcl<string> Executethetclcommandstring.TypetclmicqhelptogetalistofmICQ−specifictcl commands. SEEALSO micq(1),micqrc(5) AUTHOR ThismanpagewasoriginallycreatedbyJamesMorrison<ja2morrison@student.math.uwaterloo.ca>forareferencetoallinteractivecommandsinmICQ.Ithasbeenadaptedtocurrent usagebyRu¨ digerKuhlmann. mICQ 11